Given a binomial distribution with n = 35 and p = 0.86, would the normal distribution provide a reasonable approximation? Why or why not?

Answer by jim_thompson5910(35256)

It's not a reasonable approximation since n(1-p) = 35(1-0.86) = 4.9, which is less than 5.